The is the interest rate that a firm pays on any new debt financing. Wat after-tax cost of debt VPC) can borrow funds at an i2. An overview of a firms cost of debt The is the interest rate that a firm pays on any new debt financing. Water and Power1. The basic WACC equation The calculation of WACC involves calculating the weighted average of the required rates of return1. The basic WACC equation The calculation of WACC involves calculating the weighted average of the required rates of return
Answer #1

a)
The before tax cost of debt is the interest rate that a firm pays on any new debt financing.

b)
After tax cost of debt = Pre tax cost * (1 - tax rate)
= 10.20 * (1 - 0.25)
= 10.20 * 0.75
= 7.65%

c)
The pretax cost of debt is calculated by using the following excel formula:
=RATE(nper,pmt,pv,fv)
=RATE(15,120,-1329.55,1000)
= 8.12%
After tax cost of debt = 8.12 * (1 - 0.25)
= 8.12 * 0.75
= 6.09%

d)
rp is the symbol that represents the cost of preferred stock in the weighted average cost of capital equation.

e)
Weight on debt = 2.7 million / (2.7 million + 2.5 million + 1.8 million)
= $2.7 million / $7 million
= 0.39